OpenStudy (anonymous):

What is the equation of the line which includes points (1, 1) and (1, 5)? x = 1 y = 1 y = x + 1 y = -x - 1

OpenStudy (anonymous):

i think we can use point slope form

OpenStudy (anonymous):

would it be c?

OpenStudy (anonymous):

@abayomi12 DO YOU KNOW?

jimthompson5910 (jim_thompson5910):

Hint: Notice how the x coordinates of both points are the same

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

Your book probably has the equation for point slope form. It looks like this \[(y-y _{1})=m(x-x_{1})\] Use it to find the slope (m). Recall y=mx+b

OpenStudy (anonymous):

a

OpenStudy (precal):

plot the points, that should help you decide which one it is

OpenStudy (anonymous):

x=1

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (precal):

yw

OpenStudy (anonymous):

(1 , 1) , (1 , 5) m = (y2 - y1) / (x2 - x1) m = (5 - 1) / (1 - 1) m = 4/0 m = undefined we can see that x = 1 is in the two points, which means that x = 1 i the equation of the line.